OpenStudy (anonymous):

The Bugatti Veyron EB 16.4 Super Sport is the fastest production car in the world, with a top speed of 268 mph. It can accelerate from 0 to 100 km/hr in 2.46 seconds. What is the average acceleration of the Veyron, expressed in m/s2 , over this speed range?

OpenStudy (anonymous):

\[\frac{ V _{f}-V _{i} }{\Delta t } \] is average acceleration

OpenStudy (anonymous):

convert km/hr to m/s and solve

OpenStudy (anonymous):

so would it be 12.2 m/s?

OpenStudy (anonymous):

\[ \frac{ 100 km}{hour } * \frac{ 1000m }{ km } * \frac{ 1 hour }{ 3600 \sec }\]